OpenStudy (anonymous):

3 + [(9 ÷ 3) × 2] + [(2 × 12) ÷ 4] I wanted to make sure I worked it right I got 15

OpenStudy (vane11):

I got the same thing :)

OpenStudy (zzr0ck3r):

\(3+[(9/3)*2]+[(2*12)/4]=3+6+6 = 15 \ \ \checkmark\)

OpenStudy (anonymous):

thanks guys I have another one give me a sec to do it and I will post